OpenStudy (anonymous):

Find the zeros of the function. f(x) = 3x3 - 12x2 - 15x

Directrix (directrix):

The zeros of the function are values of x that cause the corresponding y-values to be zero. 3x^3 - 12x^2 - 15x = 0

Directrix (directrix):

There is a common factor of x in all three terms of the left side of the equation. Factor (not divide) it out. 3x^3 - 12x^2 - 15x = 0 3x ( x^2 -4 x - 5) = 0

Directrix (directrix):

The trinomial x^2 -4 x - 5 will factor over the set of Integers. 3x ( x^2 -4 x - 5) = 0 3x (x - 5) (x + 1) = 0 Use the Zero Product Property next.

Directrix (directrix):

3x (x - 5) (x + 1) = 0 3x = 0 OR x - 5 = 0 OR x + 1 = 0 Solve each of the three linear equations for x to get the three zeros. @nezzy

OpenStudy (anonymous):

0, -1, and 5